OpenStudy (anonymous):

A rock is held steady over a well and dropped. 4.6 seconds later, the sound of the rock hitting water is heard. Sound travels at a constant speed of 347 m/s. How deep is the well?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

say \(d\) is the depth of of well

ganeshie8 (ganeshie8):

say it takes \(t\) seconds for stone to hit water so it wud take \(4.6-t\) seconds for sound to travel

ganeshie8 (ganeshie8):

distance travelled by stone :- \(\large d = \frac{1}{2} \times 9.8 \times t^2\) ----------(1)

ganeshie8 (ganeshie8):

same distance is travelled by sound on the way up :- \(\large d = 347 \times (4.6-t)\) ----------(2)

ganeshie8 (ganeshie8):

two equations, two unknowns we should be able to solve them ?

OpenStudy (anonymous):

solve for t then plug into one of the above equations correct?

OpenStudy (anonymous):

to get d

ganeshie8 (ganeshie8):

yup you can do that

ganeshie8 (ganeshie8):

after all is done, you shoudl get \(d = 92.0681\)

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

thank u. i got this wrong. i will learn from it

ganeshie8 (ganeshie8):

good, np :)

ganeshie8 (ganeshie8):

did u get how i wrote the equation (1) ?

OpenStudy (anonymous):

yes i understand. u usually do not take more than one problem? is it ok if i ask u if i am stuck on sthing?

ganeshie8 (ganeshie8):

nthing like that lol, sure u can tag me... :D